Hi! I'm stuck on this question in my math and its says "Determine which of the points is a solution to the system." Hopefully I got the right topic haha :) Okay so this is the question:
y=2x-3 (4,-4);(2,1);(3,3)
4x+y=15
Thanks! -Savanah

Answer by Alan3354(69443)

"Determine which of the points is a solution to the system." Hopefully I got the right topic haha :) Okay so this is the question:
y=2x-3
(4,-4);(2,1);(3,3)
4x+y=15
-----------------
2 ways to do it.
There are 3 points, try them in both equations.
(4,-4)
y = 2x-3
-4 = 2*4 - 3 ? NG
----------
(2,1)
4x+y = 15
4*2 + 1 = 15 ? NG
--------
(3,3)
(3,3) fits both.
====================
Another way is to solve the system of equations.
y=2x-3
4x+y=15
--
Sub for y in the 2nd eqn
4x + (2x-3) = 15
6x - 3 = 15
6x = 18
x = 3
---
y = 2x-3
y = 3
--> (3,3)